During 2020, Churchill had the following convertible securities outstanding:

$200,000 of 6% convertible bonds. EAch $1000 bond is convertible into 30 shares of common stock

$100,000 of 10%, $50 par, cumulative preferred stock. Each share is convertible into 4 shares of common stock.

Churchill has an income tax rate of 30%. It's reported net income for 2020 was $210,000 and it had 26,000 shares of common stock outstanding all year.

Calculate basic and diluted earnings per share for Churchill.

Please show computations and reasoning.
Answer #1

Basic Earnings per Share = Earnings for Equity/Number of outstanding equity shares

= 210,000/26,000

= $8.077/share

Diluted:

Net Income = 210,000

Add: Interest saved on bonds net of tax = 200,000*6%(1-0.7) = 8,400

Add: Preferred Stock Dividend = 100,000*10% = 10,000

Diluted Net Income = $228,400

Diluted number of shares:

Outstanding Shares = 26,000

Convertible Bonds = 200,000*30/1000 = 6,000

Convertible Preferred Shares = 100,000*4/50 = 8,000

Diluted Number of shares = 40,000

Diluted Earnings per share = 228,400/40,000

= $5.71/share